Can someone help me solve the following problem.
3x^2 + 6 = 10

Answer by checkley77(12844)

3x^2 + 6 = 10 subtract 10
3x^2+6-10=0 combine numbers
3x^2-4=0 add 4
3x^2=4 divide by 3
x^2=4/3 take he sqrt
x=sqrt4/3 factor the fraction
x=sqrt(2*2/3)
x=2sqrt(1/3)